### v1

Abstract
Based on entropy and symmetrical uncertainty (SU), we define a metric for categorical random variables and show that this metric can be promoted into an appropriate quotient space of categorical random variables. Moreover, we also show that there is a natural commutative monoid structure in the same quotient space, which is compatible with the topology induced by the metric, in the sense that the monoid operation is continuous.
Correlation, entropy, similarity metric, distance metric, monoid
1 Introduction
Symmetric Uncertainty has been introduced in [CUP:Press_et_al:1988], and it is defined as
$SU(X,Y):=2\left[1-\frac{H(X,Y)}{H(X)+H(Y)}\right],$ (1)
where $X,Y$ are categorical random variables, $H(\cdot)$ is the entropy, and $H(\cdot,\cdot)$ is the joint entropy.
Generalization of SU for the multivariate case, as well as its properties, and, in particular, its potential to define a distance metric among categorical random variables have been explored in [Entropy:Gomez-2021, ref-Gomez2023, IS:Sosa-2018, SBMAC:Sosa-2018]. Indeed, [ref-Gomez2023] bring us to realize that entropic correlation is an example of similarity.
In this work, we further explore some properties of $SU$ in two directions. On one hand we explore some properties of the topology introduced by the metric defined in terms of $SU$ on a given space of categorical random variables. On the other hand, we explore some algebraic properties of a joint operation defined on the same space where the metric is defined. We conclude that this operation endows the metric space with a commutative monoid structure, and that it is compatible with the metric topology, in the sense that it is a continuous map.
2 Background: Entropic Correlation and Symmetric Uncertainty
In this section we recall the fundamental concepts involved in the definition of $SU$, as well as its main properties which we are going to need.
Let us consider a fixed probability space $(D,\Omega_{D},P)$, and let $X:D\to\Sigma_{X}$ be a categorical (non-numeric, nominal or cualitative) random variable with possible values $\{x_{1},\ldots,x_{k}\}$ and probability mass function $P_{X}$. The entropy $H$ of the variable $X$ is a measure of the uncertainty in predicting the value of $X$, or size of the uncertainty on $X$, and is defined as [Renyi-1961]:
$H(X):=-\sum_{i}P_{X}(x_{i})\log_{2}(P_{X}(x_{i})),$ (2)
where $P_{X}(x_{i})$ is the prior probability of the value $x_{i}$ of $X$. Any base could be used for the log, but base 2 is used for historical reasons connected to information theory [Shannon-1948]. $H(X)$ can also be interpreted as a measure of the amount of information a discrete random variable $X$ requires for describing its behavior; or more compactly, as the variety inherent to $X$.
Given another categorical random variable $Y$, the conditional entropy $H(X|Y)$ quantifies the amount of information needed to describe the outcome of $X$ given that the value of $Y$ is known. It is defined as follows:
$H(X|Y):=-\sum_{j}\left[P_{Y}(y_{j})\sum_{i}P(x_{i}|y_{j})\log_{2}(P(x_{i}|y_{j}))\right],$ (3)
where $P(x_{i}|y_{j})$ is the posterior probability of a value $x_{i}$ for the variable $X$ given that the value of variable $Y$ is $y_{j}$.
Using entropy and conditional entropy, the concept of joint entropy is introduced, which for two categorical random variables $X,Y$ is defined by
$H(X,Y)=H(Y)+H(X|Y).$ (4)
For more details, we refer the reader to [FIMU:Bouda], page 15.
Mutual Information. Using entropy and conditional entropy introduced above, the mutual information  [^1] In certain literature mutual information is denoted as $I(X;Y)$; see [ref-Fullwood2023]. $MI(X|Y)$ is introduced [MKP:Quinlan:1993]. Mutual Information measures the reduction in uncertainty about the value of $X$ when the value of $Y$ is known, and is defined as:
MI(X|Y):=H(X)H(X|Y).MI(X|Y):=H(X)-H(X|Y). (5)
Since $MI$ measures how much the information provided by $Y$ makes it easier to predict the value of $X$, it can be used as a measure of correlation. It should be noted that:
  • if $X$ and $Y$ are independent then in equation (5) $H(X|Y)=H(X)$, and hence $MI(X|Y)$ = 0. That is, under independence, knowing $Y$ gives no information about the value of $X$; and
  • if $X$ and $Y$ are fully correlated then $H(X|Y)=0$ and hence $MI(X|Y)=H(X)$ applying equation (5).
Mutual information taken as a measure of correlation, makes it possible to compare correlations with one another. For example, for any random variables $X$, $Z$ and $Y$, $MI(X|Y)>MI(Z|Y)$ means that when knowing the value of $Y$ the reduction in uncertainty about $X$ is greater than the reduction in uncertainty about $Z$, which is conducive to stating that $X$ is more correlated to $Y$ than it is $Z$.
By combining equations (4) and (5), one readily obtains
MI(X|Y)=H(X)+H(Y)H(X,Y)MI(X|Y)=H(X)+H(Y)-H(X,Y) (6)
and since $H(X,Y)$ is symmetric by direct substitution of $X,Y$ into equation (1), it results that $MI(X|Y)$ is symmetric as well, a quite convenient property for a paired measure. A formal proof can be found, for example, as Theorem 1 in [ref-Fullwood2023].
Symmetric Uncertainty. On the other hand, like with other informational association measures, MI tends to increase its value when the number of categories in $X$ and/or $Y$ increases, that is, it is biased towards high cardinality features [ML:Quinlan-1986]. Therefore, seeking to compensate for such bias, MI has to be normalized by the sum of entropies of the features. This measure, called Symmetric Uncertainty [CUP:Press_et_al:1988] is expressed as:
$SU(X,Y):=2\left[\frac{MI(X|Y)}{H(X)+H(Y)}\right]$ (7)
or, combining with equation (6),
$SU(X,Y):=2\left[1-\frac{H(X,Y)}{H(X)+H(Y)}\right].$ (8)
Revisiting the two bullet points in the previous subsection, note now that:
  • if $X$ and $Y$ are independent then, from $MI(X|Y)$ = 0, it follows that $SU(X,Y)=0$; and
  • if $X$ and $Y$ are completely correlated then from equation (5), it is obtained $MI(X|Y)=H(X)$. By symmetry of MI, $MI(Y|X)=H(Y)$ and so $SU(X,Y)=1$.
As a final observation regarding $SU$, we recall from [IS:Sosa-2018] the following results.
Lemma 2.1.
For any two categorical random variables, we have:
  1. 1)
    $0\leq SU(X,Y)\leq 1$
  2. 2)
    The Entropic Ratio, defined as
    $R(X,Y):=\frac{H(X,Y)}{H(X)+H(Y)},$ (9)
    where $H(X,Y)=H(X)+H(Y|X)$, satisfies the inequality
    $\frac{1}{2}\leq R(X,Y)\leq 1.$
It will be convenient to relate the concept of entropy of a categorical random variable to that of the entropy of a partition, as follows: let us consider again our fixed probability space $(D,\Omega_{D},P)$. A partition of $D$ is a countable collection of disjoint measurable subsets of $D$, whose union has measure $1$. We will denote by $\cal D$} the trivial partition, namely ${\cal D}=\{D\}$. Observe that we will only need finite partitions.
Given a partition $\cal Q$} of $D$, the entropy of $\cal Q$} is defined by
$H({\cal Q}):=-\sum_{Q\in{\cal Q}}P(Q)\log(P(Q)).$
[…]
$H({\cal Q}|{\cal R}):=-\sum_{Q\in{\cal Q}}\sum_{R\in\cal R}P(Q\cap R)\log\left(\frac{P(Q\cap R)}{P(R)}\right).$
Notice that, for any partition $\cal X$} of $D$ we have $H({\cal X}|{\cal D})=H({\cal X})$.
[…]
Regarding these concepts, we have the following technical results.
Lemma 2.2 (Lemma 9.1.5 in [Book:Viana-Oliveira]).
Given partitions ${\cal X},\cal{Y},{\cal Z}$, with finite entropy, the following hold:
  1. a)
    $H({\cal X}\lor{\cal Y}|{\cal Z})=H({\cal X}|{\cal Z})+H({\cal Y}|{\cal X}\lor{\cal Z})$.
  2. b)
[…]
  3. c)
    ${\cal X}\leq{\cal Y}$ if and only if $H({\cal X}|{\cal Y})=0$.
Notice that, taking ${\cal Z}={\cal D}$ in item $a)$ we get, for any two partition $\cal X$} and $\cal Y$}:
$H({\cal X\lor{\cal Y}})=H({\cal X})+H({\cal Y|X)}\implies H({\cal X})\leq H({\cal X}\lor{\cal Y})$ (10)
Also, from ${\cal Y}\leq{\cal Y}\lor{\cal Z}$ and ${\cal Z}\leq{\cal Y}\lor{\cal Z}$ we get (applying item $b)$)
$H({\cal X}|{\cal Y}\lor{\cal Z})\leq H({\cal X}|{\cal Y}),\quad H({\cal X}|{\cal Y}\lor{\cal Z})\leq H({\cal X}|{\cal Z})$ (11)
Now, if
$X:D\to\Sigma_{X}=\{x_{1},x_{2},\ldots,x_{m}\}$
is a categorical random variable defined on $D$, then it defines a partition $\cal X$} of $D$ given by the inverse images of $x_{i}$, i.e.:
${\cal X}:=\{X^{-1}(x_{i}),i=1,2,\ldots,m\}.$
The basic observation here is that the concept of entropy of a partition, and the entropy of a partition relative to another partition, recover the corresponding concepts regarding categorical random variables. In particular, given two categorical random variables $X$ and $Y$ defined on $D$, we have
$H(X,Y)=H({\cal X}\lor{\cal Y}).$
3 Topological structure of Categorical Variables
In this section we introduce a topology into a set of categorical variables. We will do so by defining a distance metric based on the SU measure. To motivate the definition, we will first explore some examples in order to get an idea of how the SU measure can give us a way to compare categorical variables.
3.1 Measuring similarity on a dataset: between observations and between variables
A set of data, or dataset, consists of $m$ observations taken across $n$ variables. Each observation is displayed as a row of $n$ individual values. Each set of values at the same position (column) constitute measurements of a variable, see Table 1.
Table 1: A generic dataset with $n$ column variables and $m$ row observations.
$X_{1}$ $X_{2}$ $X_{3}$ $X_{4}$ $X_{5}$ ... $X_{n}$
$x_{11}$ $x_{12}$ $x_{13}$ $x_{14}$ $x_{15}$ $x_{1n}$
$x_{21}$ $x_{22}$ $x_{23}$ $x_{24}$ $x_{25}$ $x_{2n}$
$x_{m1}$ $x_{m2}$ $x_{m3}$ $x_{m4}$ $x_{m5}$ $x_{mn}$
Looking row-wise at the dataset, any two observations (rows) can be compared to determine their similarity. If two rows are equal throughout all their values, we say that they are in full similarity or equivalently, we say that the distance between these observations is zero. If not all the values on the two rows are equal, several techniques are available to compute distances in this $n$-dimensional space, where a mix of data types is not unusual; see for instance Suárez-Díaz et al [UG:Suarez-2020].
We can also look at the same dataset in a column-wise manner [ref-Gomez2023]. In exploratory experiments and in model construction, one tries to establish how the value in one column (the outcome or class) results from the values in other columns, revealing a concomitance or correlation between sets of values.
Is it possible to find similarities between columns? Since each column measures a specific characteristic, and perhaps has been recorded in its own unit of measurement, we can rarely expect to find column similarity in the same way that we find similar rows. However, here one can think of similarity in a slightly different way. It is known that columns may relate to each other via correlations. Correlation does not necessarily imply causality, but it is reasonable to view high correlation between two variables $X_{i}$ and $X_{j}$ as an indicator of being linked to the same phenomenon. Because of this, we may think that if the correlation between $X_{i}$ and $X_{j}$ is high, then the two variables are “close of each other”; and conversely, low correlation between $X_{i}$ and $X_{j}$ sets these variables apart.
We address the problem of finding similarities (or from a complementary viewpoint, finding distances) between columns, variables or features of a database. That is, given any two variables of the dataset, what is the similarity between them and how can we calculate it?
For categorical variables, the Symmetrical Uncertainty measure of correlation computes this kind of similarity by comparing the joint entropy of the pair of variables with their individual entropies. So the answer to the above questions is: Yes, similarity expressed as correlation between two columns is computable as the SU. This similarity is based on variables belonging to a group of intertwined characteristics. The next example illustrates how two columns may be viewed as similar.
Example 1.
Students applying for certain internship got their scores recorded on 5 traits: Neatness, Punctuality, IQ, Attention Type, Creativity, and GotHired.
Table 2: Personality traits of students applying for an internship. Variables: Neatness (Untidy/Smooth/Refined), Creativity (Divergent/Solver/Imaginative), Punctuality (Erratic/On-time/Late), Intelligence Quotient (Low/Average/High), Attention Type (Selective/Sustained/Divided/Alternating), GotHired (No/Yes)
subjectsubject NeatnessNeatness CreativityCreativity PunctualityPunctuality IQuotientIQuotient AttentionTypeAttentionType GotHiredGotHired
1 R D L H A N
2 U S E L A N
3 R D L H A Y
4 U D E L SU Y
5 R I L H SE N
6 S S O A SU Y
7 R D L H SU Y
8 S I O A SE N
9 R S L H A N
10 S D O A A Y
11 R S L H D N
12 R I L H A N
13 U D E L SE Y
14 R D L H D Y
15 R I L H SU N
16 U I E L SU N
17 U D E L SU Y
18 S D O A A Y
19 S S O A SU N
20 R I L H D N
Table 2 presents data example for the personality traits of students applying for an internship. The $SU$ values between each variable and the gotHired class are:
  • SU(Neatness, GotHired) = 0.0535
  • SU(Creativity, GotHired) = 0.4627 (largest computed SU)
  • SU(Punctuality, GotHired) = 0.0535
  • SU(IQuotient, GotHired) = 0.0535
  • SU(AttentionType, GotHired) = 0.0192
Out of many possible predictive models, a model to predict whether a student will get hired based on a single predictor variable is perhaps the simplest one. Of all the 5 feature vs class pairs in the data, the creativity-gotHired pair has highest SU value. Let us analyze this a bit more.
Suppose we need to group our 20 cases by each category of Creativity and by each category of GotHired. When we do this, Creativity and GotHired produce about the same groupings, as seen in Table 3: creativity D is associated with GotHired=Y, and creativities S or I associate with GotHired=N.
Table 3: Counts for Creativity and Getting Hired
Creativity, Creativity, Creativity,
D S I
GotHired, Y 88 11 0
GotHired, N 11 44 66
Thus creativity is akin to getting hired, or Creativity is similar to GotHired in the correlation sense. From the point of view of a 2-way model, if we know that the value of creativity is D we can almost predict that the value of GotHired will be Y, while for creativities S or I the candidate probably won’t be hired.
The same kind of analyses could be done for the other four resulting SU values, detecting similarities with GotHired in accordance with each SU value.
Note that if we replace GotHired by its negative NotHired, all the $SU$ values would remain the same as only the probabilities of the labels (not labels themselves) participate in $SU$ computations.
3.2 Entropic equivalence and the metric structure
In many situations, given a set $\Omega$, it is useful to have a function $f\colon\Omega\times\Omega\to\mathbb{R}$ which tell us how similar or dissimilar are any given pair of elements of $\Omega$. One such a function is the well known distance metric, which is any function $d\colon\Omega\times\Omega\to\mathbb{R}$ satisfying the following conditions:
  1. 1.
    $d(x,y)\geq 0$ (non-negativity),
  2. 2.
    $d(x,y)=d(y,x)$ (symmetry),
  3. 3.
    $d(x,z)\leq d(x,y)+d(y,z)$ (triangle inequality),
  4. 4.
    $d(x,y)=0$ if and only if $x=y$. (identity of indiscernibles)
Less standardized is the concept of a similarity metric. Here we recall the definition given by Chen et al in [TCS:Chen-2009].
Definition 3.1.
Given a set $\Omega$, a real-valued function $s(x,y)$ defined on the Cartesian product $\Omega\times\Omega$ is a Similarity Metric if, for any $x,y,z\in\Omega$, it satisfies the following conditions:
  1. 1.
    $s(x,y)=s(y,x)$ (symmetry)
  2. 2.
    $s(x,x)\geq 0$ (non-negativity)
  3. 3.
    $s(x,x)\geq s(x,y)$ (self-identity)
  4. 4.
    $s(x,y)+s(y,z)\leq s(x,z)+s(y,y)$ (triangle inequality)
  5. 5.
    $s(x,x)=s(y,y)=s(x,y)$ if and only if $x=y$. (identity of indiscernibles)
In [TCS:Chen-2009], the authors also gave conditions for a similarity metric to yield a distance metric. More concretely, they established a precise relationship between normalized distances and similarity metrics. Here, normalized means that both the distance and the similarity map into $[0,1]$. For the reader convenience, let us recall their result here.
Lemma 3.2 (Corollary 1 in [TCS:Chen-2009]).
If $s(x,y)$ is a normalized similarity metric and for any $x$, $s(x,x)=1$, then $\frac{1}{2}(1-s(x,y))$ is a normalized distance metric. If, in addition, $s(x,y)\geq 0$, then 1s(x,y)1-s(x,y) is a normalized distance metric. If $d(x,y)$ is a normalized distance metric, then 1d(x,y)1-d(x,y) is a normalized similarity metric.
Looking at pairs of categorical variables, we know that the more correlated two variables are, the greater their SU value, hence SU appears to be a similarity measure. Showing that SU is a similarity metric requires the concept of indiscernibility defined for categorical random variables. Let us make this observation precise.
Definition 3.3.
Let $\cal C$ be a set of categorical random variables defined on a sample space $D$. If $X\colon D\to\Sigma_{X}$ and $Y\colon D\to\Sigma_{Y}$ are two elements in $\cal C$}, we say that they are indiscernible if they have the same histogram, modulo a permutation of the categories.
Let us recall that, given $X\colon D\to\Sigma_{X}$, the histogram is just the graph of the probability mass function $\mu_{X}\colon\Sigma_{X}\to\mathbb{R}$ given by
$P_{X}(S)=P(X^{-1}(S))=\sum_{x\in S}P(X=x).$
Thus, in the previous definition, we are asking for the graphs of $P_{X}$ and $P_{Y}$ to coincide after, if necessary, relabeling or reordering the categories in one of the sets $\Sigma_{X}$ or $\Sigma_{Y}$. In particular, for $X$ and $Y$ to be indiscernible, their codomains $\Sigma_{X}$ and $\Sigma_{Y}$ must have the same cardinality. In the following, we will refer to the indiscernibility condition as the existence of a bijection $\Sigma_{X}\to\Sigma_{Y}$ and that $P_{X}=P_{Y}$.
The indiscernibility just defined introduces an equivalence relation in $\cal C$. Let us denote by $\overline{{\cal C}}$ the space of equivalence classes. Notice that any two equivalence classes may have the same number of categories or a different number of categories.
Example 2.
Example of indiscernibles. Consider variable $X_{1}$ with values {1, 2, 3} and variable $X_{2}$ with values {A, B, C}. Suppose the following sample of 10 individuals is observed.
Table 4: Example of Indiscernibles.
RowRow $X_{1}$ $X_{2}$
1 2 B
2 3 C
3 1 A
4 1 A
5 3 C
6 3 C
7 1 A
8 3 C
9 1 A
10 3 C
We can see that the probabilities on both histograms $H(X_{1})$ and $H(X_{2})$ are 0.4, 0.1 and 0.5, so that $H(X_{1})=H(X_{2})$. Combinations of values from the two variables in the sample determine a bijection, and we have indiscernible variables with $SU(X_{1},X_{2})=1$ showing a perfect correlation.
Let us observe that the $SU$ can be promoted to the quotient space $\overline{\cal{C}}$ in the natural way, namely, if $X],[Y]\in{\cal C}$, then
$SU([X],[Y]):=SU(X,Y)$ (12)
is well defined.
Now we can state the following result, where, for simplicity, we will omit the bracket notation for equivalence classes.
Theorem 3.4.
Given a space $\mathcal{C}$ of categorical random variables, the SU as defined in equation (7), induces via (12) a positive normalized similarity metric in the space $\overline{\cal{C}}$. That is, for any $X$, $Y$ and $Z$ in $\overline{\cal{C}}$, the following conditions hold:
  1. 1.
    Symmetry: $SU(X,Y)=SU(Y,X),$
  2. 2.
    Reflexivity: $SU(X,X)\geq 0,$
  3. 3.
    Self-similarity: $SU(X,X)\geq SU(X,Y),$
  4. 4.
    Triangle inequality: $SU(X,Y)+SU(Y,Z)\leq SU(X,Z)+SU(Y,Y),$
  5. 5.
    Identity of indiscernibles: $SU(X,X)=SU(Y,Y)=SU(X,Y)$ if and only if $X=Y$.
  6. 6.
    Positivity and normality: $SU(X,Y)\in[0,1]$.
Proof.
First, note that if $X$, $Y$ and $Z$ are any categorical random variables, SU as defined in equation (7) is a composite random variable, and at the same time it is a real-valued function on the Cartesian product $\mathcal{C}\times\mathcal{C}$.
Condition 1 is straightforward by substitution into equation (7). Likewise, for the reflexivity condition 2, using equation (7) we get the self-similarity $SU(X,X)=1\geq 0$.
The self-similarity condition 3 establishes that $SU(X,X)$ is greater than or equal to any other value of SU(X, Y). This is indeed true as we have just seen that $SU(X,X)=1$ which is the maximum value achievable by the measure.
Equivalent to the triangle inequality in a distance metric, the Triangle Inequality (Condition 4) states that the similarity between $X$ and $Z$ through $Y$ is no greater than the direct similarity between $X$ and $Z$ plus the self similarity of $Y$. Let us start by rewriting inequalities for the two-variable cases, using entropic ratio notation (see Definition in Lemma 2.1). For the variables $X$, $Y$, and $Z$, we get:
$\frac{1}{2}$ $\leq R(X,Y)\leq 1,$
$\frac{1}{2}$ $\leq R(Y,Z)\leq 1,$
$\frac{1}{2}$ $\leq R(X,Z)\leq 1.$
Also note that SU(X,Y)=2(1R(X,Y))SU(X,Y)=2(1-R(X,Y)). We now show Condition 4 in four steps.
a. Working with corresponding sides, add the first two inequalities above and then subtract the third,
$1$ $\leq R(X,Y)+R(Y,Z)\leq 2$ (13)
$1/2$ $\leq R(X,Y)+R(Y,Z)-R(X,Z)\leq 1$ (14)
b. From here on use only the left and the middle members. Transpose two terms to the left, getting
$1/2-R(X,Y)-R(Y,Z)$ $\leq-R(X,Z)$ (15)
c. Add (1 + ½) to each side, then multiply each side by 2.
$1-R(X,Y)+1-R(Y,Z)$ $\leq 1-R(X,Z)+(1/2)$ (16)
$2(1-R(X,Y))+2(1-R(Y,Z))$ $\leq 2(1-R(X,Z))+2(1/2)$ (17)
d. Remembering that $SU(Y,Y)$ = 1,
$SU(X,Y)+SU(Y,Z)$ $\leq SU(X,Z)+1$ (18)
$SU(X,Y)+SU(Y,Z)$ $\leq SU(X,Z)+SU(Y,Y)$ (19)
Condition 5 results from applying $SU(X,X)$ = 1 once again, and noting that $X=Y$ is understood as equality in entropy.
Finally, notice that Condition 6 was already established in Lemma 2.1. This finalizes the proof. ∎
As a direct consequence of this Theorem, together with Lemma 3.2, we have the main result of this section.
Theorem 3.5.
Given a space $\cal C$ of categorical random variables, let $\overline{\cal C}$ be its coset defined by indiscernibility. The quantity 1SU(X,Y)1-SU(X,Y) is a normalized distance metric in $\overline{\cal C}$.
Let us notice that, since we are working with categorical variables, it may seem that the metric topology just introduced is discrete. However, this is not the case, as we show in the following theorem.
Theorem 3.6.
The metric topology introduced in $\overline{\cal C}$ is not discrete.
Proof.
Indeed, to see this, just consider two categorical random variables, $X$ and $Y$, with categories $\{x_{1},x_{2}\}$ and $\{y_{1},y_{2}\}$, and probabilities $\{p_{1},1-p_{1}\}$ and $\{q_{1}=p_{1}+\epsilon,q_{2}=1-q_{1}\}$. Then, we see that $H(X|Y)=0$, and thus
$SU(X,Y)=2\frac{H(X)}{H(X)+H(Y)}.$
Since $H(X)$ and $H(Y)$ are continuous functions of $p_{i}$ and $q_{i}$, we can make $H(X)+H(Y)\approx 2H(X)$, resulting in 1SU(X,Y)1-SU(X,Y) as small as we want by letting $\epsilon\to 0$. ∎
4 Algebraic Structure of Categorical Variables
Let $A$ and $B$, be two categorical random variables defined over a common sample space $D$. To fix the idea, let us assume that $A$ stands for income, which we consider to have three possible values, namely $\{l,m,h\}$ for “low”, “middle” and “high”, and that $B$ stands for owning or not some house, which we consider to have two possible values, namely $\{y,n\}$, for “yes” and “no”. It is quite common to produce, out of $A$ and $B$, another random variable $C$, which we will denote as C=ABC=A*B, by the following rule: Given any $p\in D$, then $C(p):=(A(p),B(p))$.
Thus, if $p\in D$ is an individual with middle income, and owns some house, then $C(p)=(m,y)$. Let us observe that the possible values for $C$ is the set
$\{(l,y),(l,n),(m,y),(m,n),(h,y),(h,n)\}.$
If we denote by $\Sigma_{A}$ the codomain of the random variable $A$ and by $\Sigma_{B}$ the codomain of the random variable $B$, then we get that the codomain of $C$ is $\Sigma_{C}=\Sigma_{A}\times\Sigma_{B}$. Let us also recall that, if $P$ is the probability measure in the sample space $D$, and $P_{A}$, $P_{B}$ are the probability mass function of $A$ and $B$, respectively, then, in the measurable product space $\Sigma_{A}\times\Sigma_{B}$ we have the joint probability mass function $P_{A}\times P_{B}\colon\Sigma_{A}\times\Sigma_{B}\to\mathbb{R}$ given by
$P_{A}\times P_{B}(x,y)=P(A=x\wedge B=y)=P(A=x|B=y)\cdot P(B=y).$
Thus, if we denote by $\cal C$} the set of all categorical random variables defined over $D$, what the previous procedure yields is an internal law of composition in $\cal C$}.
In this section we establish that this operation introduces in $\cal C$} an algebraic structure (more specifically, in the quotient space $\overline{\cal C}$ given by the indiscernibility equivalence relation), and explore some of its basic properties.
4.1 Formal definitions
Let us fix again the space $\cal C$}, and recall that it consists of all categorical random variables defined on the sample space $D$. Let us also denote again by $\overline{\cal C}$ the quotient space given by the indiscernibility equivalence relation in $\cal C$}.
Definition 4.1.
Given $A\colon D\to\Sigma_{A}$ and $B\colon D\to\Sigma_{B}$, two categorical variables in $\cal C$}. Then we define a new categorical random variable
$C\colon D\to\Sigma_{C}=\Sigma_{A}\times\Sigma_{B};C(p):=(A(p),B(p)).$
Let us denote this by C=ABC=A*B, and call it the joint of $A$ and $B$.
Proposition 4.2.
Given $A,B\in{\cal C}$, the operation
[A][B]:=[AB][A]*[B]:=[A*B]
is well defined.
Proof.
Indeed, let $A^{\prime}\colon D\to\Sigma_{A^{\prime}}$ and $B^{\prime}\colon D\to\Sigma_{B^{\prime}}$ be two categorical random variables equivalent to $A$ and $B$, respectively. Then, the bijections $\Sigma_{A^{\prime}}\to\Sigma_{A}$ and $\Sigma_{B^{\prime}}\to\Sigma_{B}$ yield a bijection $\Sigma_{A^{\prime}}\times\Sigma_{B^{\prime}}\to\Sigma_{A}\times\Sigma_{B}$. Also, since $A\sim A^{\prime}$, we have $P_{A}=P_{A^{\prime}}$ and similarly, $P_{B}=P_{B^{\prime}}$. It follows that the joint probability measure $P_{A^{\prime}*B^{\prime}}$ and $P_{A*B}$ are also equal. Thus, we get that $A^{\prime}*B^{\prime}$ and ABA*B are equivalent. ∎
Proposition 4.3.
The joint operation $\colon\overline{\cal C}\times\overline{\cal C}\to\overline{\cal C}$ is associative and commutative.
Proof.
Consider $A],[B],[C]\in\overline{\cal C}$, and let us pick the representatives
$A\colon D\to\Sigma_{A};\quad B\colon D\to\Sigma_{B};\quad C\colon D\to\Sigma_{C}.$
Then, we have $B*C:D\to\Sigma_{B}\times\Sigma_{C}$, as well as $A*B\colon D\to\Sigma_{A}\times\Sigma_{B}$. Then, we get
$A*(B*C)\colon D\to\Sigma_{A}\times(\Sigma_{B}\times\Sigma_{C})=\Sigma_{A}\times\Sigma_{B}\times\Sigma_{C}$
and also
$(A*B)*C\colon D\to(\Sigma_{A}\times\Sigma_{B})\times\Sigma_{C}=\Sigma_{A}\times\Sigma_{B}\times\Sigma_{C}.$
Finally, we also have
$\begin{split}[A*(B*C)](p)&=(A(p),(B*C)(p))=(A(p),(B(p),C(p)))\\ &=((A(p),B(p)),C(p))=[(A*B)*C](p).\end{split}$ (20)
Hence, we get A(BC)=(AB)CA*(B*C)=(A*B)*C. Passing to equivalence classes, we get the associativity
[A]([BC])=([A][B])[C].[A]*([B*C])=([A]*[B])*[C].
For the commutativity, let us observe that $B*A\colon D\to\Sigma_{B}\times\Sigma_{A}$. Hence, we have a bijection between $\Sigma_{A*B}$ and $\Sigma_{B*A}$. Also, for the probability mass function we have
$P_{A*B}(x,y)=P(A=x\wedge B=y)=P(B=y\wedge A=x)=P_{B*A}(y,x).$
Thus, we get that ABA*B and BAB*A are equivalent. Hence
[A][B]=[AB]=[BA]=[B][A],[A]*[B]=[A*B]=[B*A]=[B]*[A],
as claimed. ∎
So far, we have an internal operation in the space $\overline{\cal C}$, which is associative and commutative. Next we identify a categorical variable (or better, an equivalence class of categorical variables) that plays the role of neutral element.
Proposition 4.4.
Let $\Phi\colon D\to\{\phi\}$ be a categorical random variable whose only possible outcome is the singleton $\{\phi\}$. Then, for any other random variable $A\colon D\to\Sigma_{A}$, we have
$A]*[\Phi]=[A].$
Proof.
Let us first observe that any random variable on $D$ having a singleton as the only possible outcome is equivalent to $\Phi$.
Now we observe that $\Sigma_{A*\Phi}=\Sigma_{A}\times\{\phi\}$, so we have a bijection $\Sigma_{A}\to\Sigma_{A*\Phi}$ given by $x\mapsto(x,\phi)$. Also, the probability mass function of $\Phi$ is $P_{\Phi}(\phi)=1$, thus we get
$P_{A*\Phi}(x,\phi)=P(A=x|\Phi=\phi)\cdot P(\Phi=\phi)=P(A=x)=P_{A}(x).$
Hence, $A*\Phi$ and $A$ have the same histogram.
Thus, we conclude that $A*\Phi$ is equivalent to $A$. It follows that
$A]*[\Phi]=[A*\Phi]=[A],$
as claimed. ∎
Let us summarize the previous discussion in the following theorem
Theorem 4.5.
The joint operation on $\overline{{\cal C}}$ yields a commutative monoid structure.
4.2 Compatibility of the algebraic and topological structures
Let us recall that, in the previous section, we found a metric topology structure in the set $\overline{{\cal C}}$, given by the metric d=1SUd=1-SU. Now we will show that the algebraic structure given on the same set by the joint operation and the topological structure in $\overline{\cal C}$} are compatible.
To do so, let us first rewrite the distance function on a more symmetric form. For that, we recall the symmetry of mutual information, namely:
M(X|Y)=H(X)H(X|Y)=H(Y)H(Y|X)=M(Y|X).M(X|Y)=H(X)-H(X|Y)=H(Y)-H(Y|X)=M(Y|X).
Hence, for the distance d=1SUd=1-SU we get
$\begin{split}d(X,Y)=1-SU(X,Y)&=1-\frac{2MI(X|I)}{H(X)+X(Y)}\\ &=\frac{H(X)+H(Y)-H(X)+H(X|Y)-H(Y)+H(Y|X)}{H(X)+H(Y)}\\ &=\frac{H(X|Y)+H(Y|X)}{H(X)+H(Y).}\end{split}$ (21)
Also, let us observe that, given categorical random variables $X\colon D\to\Sigma_{X}$ and $Y\colon D\to\Sigma_{Y}$, we have
$H(X,Y)=H(X*Y)=H({\cal X}\lor{\cal Y}),$
where $\cal X$} and $\cal Y$} are the partitions of $D$ associated to the variables $A$ and $Y$, respectively.
Now we have all the tools in place to prove the aforementioned compatibility.
Theorem 4.6.
The joint operation
$\colon\overline{{\cal C}}\times\overline{{\cal C}}\to\overline{{\cal C}}$
is continuous with respect to the topology on $\overline{{\cal C}}$} given by d=1SUd=1-SU and the natural product topology induced in $\overline{{\cal C}}\times\overline{{\cal C}}$.
Proof.
Let $X,Y,Z,W$ be categorical random variables on $D$, for simplicity, we will avoid the equivalence class notation, and we will also use the same symbols for their associated partitions. Then we have
$d(X*Y,Z*W)=\frac{H(X*Y|Z*W)+H(Z*W|X*Y)}{H(X*Y)+H(Z*W)}$ (22)
Applying item $a)$ from Lemma 2.2, and the inequality given in equation  (11) to the terms in the numerator, we get:
$\begin{split}H(X*Y|Z*W)&=H(X|Z*W)+H(Y|X*Z*W)\leq H(X|Z)+H(Y|W)\\ H(Z*W|X*Y)&=H(Z|X*Y)+H(W|Z*X*Y)\leq H(Z|X)+H(W|Y).\end{split}$ (23)
Thus, we get
$d(X*Y,Z*W)\leq\frac{H(X|Z)+H(Z|X)}{H(X*Y)+H(Z*W)}+\frac{H(Y|W)+H(W|Y)}{H(X*Y)+H(Z*W)}$ (24)
Now we apply inequality  (10) to each summand in the denominators, to get
$\begin{split}d(X*Y,Z*W)&\leq\frac{H(X|Z)+H(Z|X)}{H(X)+H(Z)}+\frac{H(Y|W)+H(W|Y)}{H(Y)+H(W)}\\ &=d(X,Z)+d(Y,W)\end{split}$ (25)
Thus, the mapping is contractive, hence, continuous. ∎
5 Concluding Remarks
In this paper we have demonstrated that 1SU1-SU, where $SU$ is the entropic correlation between categorical random variables, is a distance metric, thus endowing the space of categorical random variables with a topological structure. We have also showed that a natural joint operation (denoted *) endows the same space with an algebraic structure, concretely, a commutative monoid structure. Finally, we have proven that the topological and algebraic structures are compatible, given that the joint operation is continuous with respect to the topological structure.
The compatibility provides a smooth and straightforward way for the user to interpret and understand the value output from SU. This way, SU has gained intuitivity in the view of the statistics practitioner, giving everyone the ability to operate with and employ this entropic (non-parametric) correlation in much the same way as the Pearson (parametric) correlation has been employed throughout the years.
Equipped with compatible topology and algebraic structure, $SU$ further formalizes CRVs as repositories and proper representations of the concepts conveyed by the variable’s categories. Thus, qualitative variables not only can be counted in their categories, but their mutual distances now tell us about similarities or likely associations. These new possibilities and measures are offered in an environment of mathematical rigor.
The authors are now working to extend the findings to the MSU measure on $n$ categorical variables. Note that the MSU measure could as well be called Multivariable Entropic Correlation, as uncertainties get mutually cancelled when computing correlation.
Acknowledgments
The authors acknowledge Professor Mauricio Poletti for fruitful discussions on the topics of this paper. The authors acknowledge support from FEEI-CONACYT-PROCIENCIA-PRONII and the ARASY project, number ESTR01-23. CES is also affiliated with the Centro de Investigación en Matemática - CIMA, Asunción, Paraguay.